Problem 2. [05 Marks] Ali & Sons, Inc., purchased a new machine on Sept 1, 2020 at a cost of Rs.180,000. The machine’s estimated useful life at the time of the purchase was five years, and its residual value was Rs.10,000. Required: Prepare a complete depreciation schedule, beginning with calendar year 2020, under each of the methods listed below (assuming half-year convention): 1. 200 percent declining-balance. 2. 150% declining-balance, switching to straight-line when that maximizes the expense


1
Expert's answer
2021-06-21T12:02:29-0400
"\frac{180000-10000}{5}=34000"

1. For 2020

"\frac{4}{12}\times34000\times2=22667"

For 2021


"\frac {180000-10000-22667}{5}\times 2=66933"

for 2022


"\frac{147333-66933}{5}\times2=32160"

for 2023


"\frac{80400-32160}{5}\times2=19296"

for 2024


"\frac{48240-19296}{5}\times2=11585"

for 2025


"\frac{28944-11585}{5}\times2=6944"

2.

for 2020


"170000-\frac{4}{12}\times0.15\times180000=161000"

for 2021


"161000\times(1-0.15)=136850"

for 2022


"136850\times(1-0.15)=116323"

for 2023


"116323\times(1-0.15)=98874"

for 2024


"98874\times(1-0.15)=84043"

for 2025


"84043\times(1-0.15)=71437"